# Inverse Reinforcement Learning with Switching Rewards and History Dependency for Characterizing Animal Behaviors

### v1

@@ B.2 Empirical runtime
Our SWIRL implementation leverages the advantages of JAX, including just-in-time (JIT) compilation and vectorization, to achieve high computational efficiency. For S-2 experiments across all three datasets (gridworld, labyrinth, and spontaneous behavior), SWIRL converges within 15–30 minutes on a V100 GPU, which takes 50–100 EM iterations. For longer $L$, a S-4 experiment on labyrinth with 50 EM iterations take 2-3 hours to finish on a L40S GPU. We switch to a L40S GPU for L=4 due to the V100 GPU’s insufficient memory capacity.
B.3 Discussion on discount factor $\gamma$ and temperature $\alpha$
We set the discount factor $\gamma=0.95$, a standard choice in RL and IRL literature. For the mouse spontaneous behavior dataset, we also tested a smaller $\gamma=0.7$, as previous literature (Markowitz et al., 2023) suggested this value as optimal for the dataset. However, we observed that the results learned by SWIRL were highly similar for both discount factors, indicating that the choice of $\gamma$ had minimal impact on performance in this case.
@@ B.3 Discussion on discount factor $\gamma$ and temperature $\alpha$
B.4 Discussion on the number of hidden modes $Z$ and history length $L$
@@
B.4.1 $Z$ in labyrinth experiment
Figure 5: Water-restricted labyrinth experiment with different number of hidden modes $Z$. (A) Held-out test LL for each model across different number of hidden modes. The shaded area represents the total area that falls between one standard deviation above and below the mean. (B) Segments of a trajectory from held-out test data, predicted by SWIRL (S-2) with $Z=3$ and $Z=4$. The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (C) Inferred reward maps from SWIRL (S-2) with $Z=4$: water, home, and two explore maps. To enhance visualization, the inferred reward $r(s_{t},s_{t-1})$ was averaged over $s_{t-1}$ to produce $r(s_{t})$.
@@ B.4.1 $Z$ in labyrinth experiment
B.4.2 $L$ in labyrinth experiment
Figure 6: Water-restricted labyrinth experiment with different action-level history length $L$. (A) Boxplot showing held-out test LL, with the x-axis representing the four different models from $L=1$ to $L=4$. Outlier selection method is described in Appendix B.6. (B) Segments of a trajectory from held-out test data, predicted by the four SWIRL models. The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (C) Inferred reward maps from SWIRL (S-4): water, home, and explore. To enhance visualization, the inferred reward $r(s_{t},s_{t-1},s_{t-2},s_{t-3})$ was averaged over $(s_{t-1},s_{t-2},s_{t-3})$ to produce $r(s_{t})$.
@@ B.4.2 $L$ in labyrinth experiment
B.4.3 $Z$ and $L$ in mouse spontaneous behavior experiment
@@ B.5 Robustness of SWIRL
B.6 Outlier selection in box plot
@@ C.1 An example labyrinth trajectory
To further explore the hidden mode segments of the trajectory from held-out test data presented in Fig. 3F, we visualized the segments corresponding to each hidden mode in this trajectory in detail (Fig. 8B). The visualization reveals that ”home” segments can remain valid even without a visit to state 0, as these segments often instead terminate at state 1 or state 2, which are next to state 0.

### v2

@@ B.2 Empirical runtime
Our SWIRL implementation leverages the advantages of JAX, including just-in-time (JIT) compilation and vectorization, to achieve high computational efficiency. For S-2 experiments across all three datasets (gridworld, labyrinth, and spontaneous behavior), SWIRL converges within 15–30 minutes on a V100 GPU, which takes 50–100 EM iterations. For longer $L$, a S-4 experiment on labyrinth with 50 EM iterations take 2-3 hours to finish on a L40S GPU.
B.3 Evaluation metrics
The test log-likelihood is computed from the forward variable $\alpha$, where $\log p(s_{1:T},a_{1:T})=\log\sum_{z}\alpha_{T,z}$. Detail of the forward variable $\alpha$ computation is described in Appendix A.2. For test hidden mode segmentation accuracy, we use the Viterbi algorithm to predict the hidden mode of each timepoint and compute the percentage of correct hidden mode prediction compared to ground truth data. For reward correlation, we compute the Pearson correlation between learnt reward and ground truth reward or dopamine fluctuation levels.
B.4 Environment setup
5x5 Gridworld: State space: Discrete(25); Action space: Discrete(5). The 5 actions are ‘up’, ‘left’, ‘down’, ‘right’, ‘stay’.
Labyrinth: State space: Discrete(127); Action space: Discrete(4). The labyrinth has a binary-tree structure and the 4 actions are ‘move in left’, ‘move in right’, ‘move out left’, ‘move out right’.
Spontaneous Behavior: State space: Discrete(9); Action space: Discrete(9). Each state is a MoSeq syllable that defines a behavior motif. Each action represents a next state that the agent transfers into.
B.5 Discussion on discount factor $\gamma$ and temperature $\alpha$
We set the discount factor $\gamma=0.95$, a standard choice in RL and IRL literature. For the mouse spontaneous behavior dataset, we also tested a smaller $\gamma=0.7$, as previous literature Markowitz et al. (2023) suggested this value as optimal for the dataset. However, we observed that the results learned by SWIRL were highly similar for both discount factors, indicating that the choice of $\gamma$ had minimal impact on performance in this case.
@@ B.5 Discussion on discount factor $\gamma$ and temperature $\alpha$
B.6 Discussion on the number of hidden modes $Z$ and history length $L$
@@
B.6.1 $Z$ in labyrinth experiment
Figure 5: Water-restricted labyrinth experiment with different number of hidden modes $Z$. (A) Held-out test LL for each model across different number of hidden modes over 10 runs. The shaded area represents the total area that falls between one standard deviation above and below the mean. (B) Segments of a trajectory from held-out test data, predicted by SWIRL (S-2) with $Z=3$ and $Z=4$. The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (C) Inferred reward maps from SWIRL (S-2) with $Z=4$: water, home, and two explore maps. To enhance visualization, the inferred reward $r(s_{t},s_{t-1})$ was averaged over $s_{t-1}$ to produce $r(s_{t})$.
@@ B.6.1 $Z$ in labyrinth experiment
B.6.2 $L$ in labyrinth experiment
Figure 6: Water-restricted labyrinth experiment with different action-level history length $L$. (A) Boxplot showing held-out test LL over 10 runs, with the x-axis representing the four different models from $L=1$ to $L=4$. Outlier selection method is described in Appendix B.8. (B) Segments of a trajectory from held-out test data, predicted by the four SWIRL models. The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (C) Inferred reward maps from SWIRL (S-4): water, home, and explore. To enhance visualization, the inferred reward $r(s_{t},s_{t-1},s_{t-2},s_{t-3})$ was averaged over $(s_{t-1},s_{t-2},s_{t-3})$ to produce $r(s_{t})$.
@@ B.6.2 $L$ in labyrinth experiment
B.6.3 $Z$ and $L$ in mouse spontaneous behavior experiment
@@ B.7 Robustness of SWIRL
B.8 Outlier selection in box plot
@@ C.1 An example labyrinth trajectory
To further explore the hidden mode segments of the trajectory from held-out test data presented in Fig. 3E, we visualized the segments corresponding to each hidden mode in this trajectory in detail (Fig. 8B). The visualization reveals that ”home” segments can remain valid even without a visit to state 0, as these segments often instead terminate at state 1 or state 2, which are next to state 0.
Appendix D Appendix D
D.1 Model-free SWIRL on a simulated gridworld environment
Figure 9: True and discovered reward maps in model-free SWIRL (S-2) simulation experiment. The color scale represents reward values ranging from 0 to 1. The home reward is defined as $r(s_{t})$, while the water reward depends on both the current and previous locations, $r(s_{t},s_{t-1})$. To present the water reward, each location is divided into five groups, as detailed in the table on the far right. For example, the polygon in the first row represents the reward value when $s_{t}$ is the current location and $s_{t-1}$ is the location to the right. Light grey indicates an impossible transition where no reward exists.
To demonstrate that SWIRL also supports model-free IRL, we replaced the model-based soft-Q iteration IRL approach used in main experiments with a model-free IRL method IQ-Learn (Garg et al., 2021). Garg et al. (2021) shows that we can convert the MaxEnt IRL objective $L(\pi,r)$ with a convex reward regularizer to an offline inverse Q-learning objective:
[…]
with $V(s)=\log\sum_{a}\exp{Q(s,a)}$. $\phi$ is the concave function from a convex conjugate. When using $\chi^{2}$-divergence, $\phi(x)=x-\frac{x^{2}}{4}$. The reward can be obtained as $r(s,a)=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[\log\sum_{a^{\prime}}\exp{Q\left(s^{\prime},a^{\prime}\right)}]$. By marginalizing over actions, one can similarly obtain a state-only reward $r(s)$.
Since optimizing the reward by maximizing $\log\pi$ corresponds to the Lagrangian dual of the standard MaxEnt IRL approach (Zeng et al., 2022), we can maximize the IQ-Learn objective $J(Q)$ instead of the $\log\pi$ term in SWIRL EM auxiliary function (Eq. 3). Therefore, for a trajectory $\xi$, SWIRL can perform IRL with the following objective:
[…]
with $V_{z_{t}}(s^{L})=\log\sum_{a}\exp{Q_{z_{t}}(s^{L},a)}$.
We tested this IQ-Learn-based, model-free SWIRL (S-2) on simulated trajectories in the same 5x5 gridworld environment as Sec. 4.1. The recovered reward maps achieved a Pearson correlation of 0.737 relative to the true reward maps. Although the model-free SWIRL recovered reward maps (Fig. 9) appear noisier than those from the model-based approach in Sec. 4.1 (Fig. 2A), they still accurately capture the critical high-reward regions.